A manufacturer calculates that the price elasticity of demand for his /her good is 2.2 .it is
decided to lower the price of the good from birr 6 to 5. At this lower price sales (TR) increases
to 15 birr per week. Calculate the quantity sold per week at the original price.
1
Expert's answer
2021-08-18T07:03:19-0400

"TR=P\u0445Q_d"

"Q_d=\frac{15}{5}=3" - final quantity with lower price.

"E_P=\frac{\%\Delta Q}{\%\Delta P}"

"\%\Delta P=\frac{5-6}{6}=0.1667"

"\frac{\%\Delta Q}{0.1667}=2.2" , then "\%\Delta Q=2.2*0.1667=0.3667"

"\frac{3-Q_0}{Q_0}=0.3667"

"Q_0=2.195" -the initial quantity sold per week.